a boat can travel 9 miles up a river in the same ammount of time it takes to travel 11 miles down the same river. if the current is 2 miles per hour, what is the speed of the boat in still water?
Answer by checkley75(3666)

D=RT OR DISTANCE =RATE*TIME
9=(R-2)T UPSTREAM
11=(R+2)T DOWNSTREAM
USING THE FIRST EQUATION SOLVING FOR T WE GET
T=9/(R-2)
SOLVING FOR T IN THE SECOND EQUATION WE GET
T=11/(R+2)
NOW SETTING THESE 2 EQUATION EQUAL WE GET
9/(R-2)=11/(R+2) NOW WE CROSS MULTIPLY & WE GET:
11(R-2)=9(R+2)
11R-22=9R+18
11R-9R=18+22
2R=40
R=40/2
R=20 WHICH IS THE RATE OF THE BOAT IN STILL WATER.
PROOF
9/18=11/22
1/2=1/2